OpenStudy (samwisegamge):

which point represents a solution to the equation x > -2 y >or equal to 2x + 7

OpenStudy (anonymous):

x-4y+z =6 2x+5y-z=7 2x-y-z = Z=6/ y=/ x=4

OpenStudy (samwisegamge):

where did you get z from